Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 232     Entry time: 08 Jun 2023, 11:02
Author: Johannes Moeller 
Group: MID 
Subject: Issue 

Addition to previous issue:

Last incident (of overall 4), we had 0 bunches, moved in XTD1 ATT 0.6 and 1.2 mm CVD diamond in and 2.4mm CVD out. However, vacuum started to become bad already two days ago.

Update from vacuum OCD: Seems to be related to one leaky below of 2.4 mm CVD arm. Operational consequences are discussed right now.

Attachment 1: Screenshot_from_2023-06-08_11-03-50.png  51 kB  Uploaded 08 Jun 2023, 12:04  | Hide | Hide all
Screenshot_from_2023-06-08_11-03-50.png
Attachment 2: Screenshot_from_2023-06-08_11-08-06.png  44 kB  Uploaded 08 Jun 2023, 12:08  | Hide | Hide all
Screenshot_from_2023-06-08_11-08-06.png
ELOG V3.1.4-7c3fd00